Circular No. 8308

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003lw AND GRB 031203 G. Tagliaferri, S. Covino, D. Fugazza, Istituto Nazionale di Astrofisica

(INAF), Brera; G. Chincarini, Milano-Bicocca University and INAF, Brera; D. Malesani, Scuola Internazionale Superiore di Studi Avanzati, Trieste; M. Della Valle, INAF, Arcetri; and L. Stella, INAF, Rome, on behalf of a larger collaboration, report on optical and near-infrared sub-arcsecond observations, made with the European Southern Observatory’s Very Large Telescope (VLT) at Paranal and New Technology Telescope (+ SofI) at La Silla, of the apparent host galaxy of GRB 031203 (http://gcn.gsfc.nasa.gov/ gcn/gcn3/2459.gcn3; GCN 2475), made up to 3 months after the discovery of GRB 031203. The object is located at α ≃ 8^{h}02^{m}30^{s}.2, δ ≃ −39^{o}51′03′′ (equinox 2000.0; GCN 2473, 2490). Preliminary analysis of the R- and I-band photometric data shows a variability fully consistent with the evolution of a supernova, similar to that of SN 1998bw (Galama et al. 1998, Nature 395, 670) at z = 0.105 (GCN 2482), whose maximum occurred ∼ 20 (observed) days after the γ-ray outburst; R magnitudes (typical relative errors 0.02 mag, plus 0.05-mag calibration uncertainty): 2003 Dec. 15.31 UT, 20.23; 17.28, 20.22; 23.3, 20.17; 28.3, 20.18; 30.24, 20.19; 2004 Jan. 16.19, 20.37; Feb. 20.10, 20.48. This may suggest that the γ-ray burst and the supernova went off almost simultaneously (± 4 days). Three FORS1 and FORS2 VLT spectra (range 380–750 nm; resolution ∼ 0.5 nm); were taken on 2003 Dec. 20, 30, and 2004 Mar. 1. After subtracting the contribution of the host galaxy from the spectra obtained on Dec. 20.31 and 30.27 (1.5 hr each), broad bumps were detected at ≈ 540 nm and 650 nm, similar to those observed in the spectrum of the peculiar type-Ic supernova 1998bw near maximum light (Patat et al. 2001, Ap.J. 555, 900), leading to the conclusion that the observed rebrightening is due to a supernova component, hereby designated as SN 2003lw.

SUPERNOVAE 2004at AND 2004au Corrigendum. Matheson et al. erroneously reported their spectroscopic

information concerning SN 2004au (cf. IAUC 8305) under the heading “SN 2004at” (thus the erroneous labelling on IAUC 8304). They add that a spectrum of SN 2004at (cf. IAUC 8302), obtained by P. Berlind on Mar. 19.24 UT, shows it to be a type-Ia supernova a few days before maximum; adopting the NED recession velocity of 6935 km/s for the host galaxy, the supernova expansion velocity is 12700 km/s for Si II (rest 635.5 nm).

2004 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green